# Everybody Prune Now: Structured Pruning of LLMs with Only Forward Passes

### v3

@@ Appendix I How many perturbative samples are reasonable?
Using $\mathrm{ns}_{\mathrm{sub-models}}=50$ results in an model with NaN perplexity on the Wikitext validation set. We posit that this is because of the LLaMA models are half precision, and removing the wrong modules can result in activations going outside of the FP16 dynamic range for unique data-points. Note that we are able to recover good performance of the model after fine-tuning though (we do not observe NaNs with the Wikitext-2 training data). This indicates that Bonsai actually recovers good modules even using as few samples as 50 sub-models.
Appendix J Mistral-7B Experiment Details
@@
The Mistral-7B model architecture differs from the LLaMA architecture in its use of group query attention and sliding window attention in lieu of the standard self-attention used in most transformer-based models like LLaMA Jiang et al. (2023). We factor these differences into consideration in the implementation of Bonsai for Mistral. For the experiments that produced the results below, all Bonsai hyper-parameters are the same as Appendix C.3.
Table 15 presents the test perplexity results for Mistral-7B under different pruning methods. Considering the fully-structured sparsity nature of Bonsai, it achieves a test perplexity of 47.5 without post-pruning adaptation, with 1.66$\times$ inference speedup. After performing post-pruning adaptation on our pruned Mistral-7B, perplexity dropped drastically to 10.08. Note that the reported results of Wanda-pruned Mistral-7B are not fine-tuned afterward; if they were, their results would be marginally better than Bonsai’s results. However, as shown in Table 2, latency speedup would have dropped rapidly, while Bonsai stays the same at 1.66$\times$.
@@
We further investigate the pruning habits of Bonsai by examining the pruned layers of Mistral, as shown in Figure 6. We notice a recurring theme: when an attention layer is significantly altered, it leads to compensation in the next layers within the sequence. This adaptive behavior, termed the ”Hydra effect” by (McGrath et al., 2023), implies that the layers within a language model interact in a way that changes in one layer prompt adjustments in another. (McGrath et al., 2023) specifically mentioned that when one attention layer was removed from a language model, the model was still able to self-repair and produce similar outputs; but it did so by relying more heavily on other layers.
Figure 6: Mistral’s pruned attention layers. The heavily pruned layers are usually preceded by or sandwiched between lightly-pruned layers, exhibiting the self-repairing ”Hydra effect” McGrath et al. (2023).
Should Bonsai prune iteratively? Table 16 demonstrates the benefits of using Bonsai in an iterative fashion. Pruning slowly ($p_{\mathrm{iter}}=0.05$) yields the best results, but this comes at the cost of increasing the total time to prune the model. The performance gap between values of $p_{\mathrm{iter}}$ persists even after post-pruning adaptation, indicating that slower pruning allows for more accurate estimates of module importance.
Table 16: Varying $p_{\mathrm{iter}}$. Wikitext-2 perplexity of LLaMA-2 7B pruned to 50% sparsity. See Appendix E for experiment details.

### v4

@@ Appendix F How many perturbative samples are reasonable?
Using $\mathrm{ns}_{\mathrm{sub-models}}=50$ results in a model with NaN perplexity on the Wikitext validation set. We posit that this is because of the LLaMA models are in half precision, and removing the wrong modules can result in activations going outside of the FP16 dynamic range for unique data points. Note that we are able to recover good performance of the model after fine-tuning though (we do not observe NaNs with the Wikitext-2 training data). This indicates that Bonsai actually recovers good modules even using as few samples as 50 sub-models.
Appendix G LlaMA-3-8B Experiments
In our efforts toward generalization, we also validate Bonsai on LLaMA-3-8B and use the same experimental setup as our LLaMA-2-7B experiments (hyperparameters in Appendix A.2) and prune to 50% sparsity on Wikitext-2.
Table 16: LLaMA-3-8B at 50% sparsity on Wikitext-2
Method Wikitext-2 PPL
LLaMA-3-8B (base) 6.14
Bonsai (w/o PPA) 21.36
Bonsai (w/ PPA) 10.37
The results show comparable trends to LLaMA-2-7B in the main text: Bonsai operates within accessible memory budgets while achieving competitive accuracy after post-pruning adaptation. This confirms that Bonsai’s approach generalizes across LLaMA model versions.
Appendix H Mistral-7B Experiment Details
@@
The Mistral-7B model architecture differs from the LLaMA architecture in its use of group query attention and sliding window attention in lieu of the standard self-attention used in most transformer-based models like LLaMA Jiang et al. (2023). We factor these differences into consideration in the implementation of Bonsai for Mistral. For the experiments that produced the results below, all Bonsai hyper-parameters are the same as Appendix A.2.
[…]
@@
We further investigate the pruning habits of Bonsai by examining the pruned layers of Mistral, as shown in Figure 5. We notice a recurring theme: when an attention layer is significantly altered, it leads to compensation in the next layers within the sequence. This adaptive behavior, termed the "Hydra effect" by (McGrath et al., 2023), implies that the layers within a language model interact in a way that changes in one layer prompt adjustments in another. (McGrath et al., 2023) specifically mentioned that when one attention layer was removed from a language model, the model was still able to self-repair and produce similar outputs; but it did so by relying more heavily on other layers.
Figure 5: Mistral’s pruned attention layers. The heavily pruned layers are usually preceded by or sandwiched between lightly-pruned layers, exhibiting the self-repairing "Hydra effect" McGrath et al. (2023).
Appendix I Should Bonsai prune iteratively?
Table 18 demonstrates the benefits of using Bonsai in an iterative fashion. Pruning slowly ($p_{\mathrm{iter}}=0.05$) yields the best results, but this comes at the cost of increasing the total time to prune the model. The performance gap between values of $p_{\mathrm{iter}}$ persists even after post-pruning adaptation, indicating that slower pruning allows for more accurate estimates of module importance.
Table 18: Varying $p_{\mathrm{iter}}$. Wikitext-2 perplexity of LLaMA-2 7B pruned to 50% sparsity. See Appendix C for experiment details.